OpenStudy (hockeychick23):

Suppose a bag contains 8 silver coins and 2 bronze coins. What is the probability of randomly choosing a silver coin, not replacing it, and then randomly choosing another silver coin?

OpenStudy (anonymous):

okay, your system has 2 steps First: 8 silver coins, 2 bronze coins So the prob. of getting a silver coin is 8/10=4/5 Second: 7 silver coins, 2 bronze coins So the prob. is 7/9 The prob. of both things would be P=P1*P2 Prob=(4/5)*(7/9)